OpenStudy (anonymous):

Find the point on the plane 2x -3y +z = 3 that is closest to the origin so I picked the point P(1,0,1). then I use a projection? and something else?

OpenStudy (turingtest):

I can only do this problem with lagrange multipliers, which is a multi variable calc tehnique

OpenStudy (anonymous):

too bad I think Im getting it now.. Im in linear algebra.

OpenStudy (anonymous):

It can be done using gradient Find the square distance from (0,0,0} to {x,y,3-2x +3y Minimize it (Solve for gradient ={0,0,0} You are done

OpenStudy (anonymous):

You end up needing to minimize the function \[ f(x,y)=5 x^2-12 x y-12 x+10 y^2+18 y+9 \]

OpenStudy (anonymous):

I got (3/7,-9/14,3/14) does this look correct?

OpenStudy (turingtest):

@eliassaab won't that require multi variable calc techniques like lagrange multipliers?

OpenStudy (anonymous):

\[ gadienr(f)={-12 + 10 x - 12 y, 18 - 12 x + 20 y, 0} \] You find where the gradient is zero at \[ \left\{\left\{x\to \frac{3}{7},y\to -\frac{9}{14}\right\}\right\} \]

OpenStudy (anonymous):

Yes, it is correct

OpenStudy (anonymous):

\[ grad(f)=\{10 x-12 y-12,-12 x+20 y+18,0\} \]

OpenStudy (anonymous):

great thank you!. and the distance would be the sqrt((x2-x1)^2 +(y2-y1)^2 + (z2-z1)^2)?

OpenStudy (anonymous):

It only requires the use of the gradient. Lagrange Multipliers does use gradient technique. You can use it in this problem if you want, but you do not have to. Of course, we are using multi-variable calculus/

OpenStudy (anonymous):

It is passed my bedtime, I am going to sleep.